in the equation 9=-1/4x^2+x, what does x equal?
Answer by lwsshak3(11628)

in the equation 9=-1/4x^2+x, what does x equal?
9=-1/4x^2+x
x^2/4-x+9=0
x^2-4x+36=0
a=1, b=-4, c=36
discriminant=(b^2-4ac)=(16-4*1*36)<0
no real solution